OpenStudy (anonymous):

assume Y varies inversely with X, so if Y=11 when X is 20, find X when Y is 66, round to the nearest tenth

OpenStudy (anonymous):

@E.ali help!

OpenStudy (anonymous):

If y varies inversely with x, and y is six times bigger, that means x needs to be six times smaller. It would seem x= 10/3 is about right.

OpenStudy (anonymous):

x_________________20_________________? y__________________11__________________66 ?=66/11=6 6*20=120 x=120

OpenStudy (strawberry17):

y 11 ? -- = --- --- x 20 66 I would multiply 11 by 66 (=726) and then divide that by 20 to get x :)

OpenStudy (strawberry17):

how is it we all got different answerss... @E.ali

OpenStudy (anonymous):

Because you have a fall to say that :) You said : 11/20 but its 11/66 :) Thank you !:)

OpenStudy (strawberry17):

lol

OpenStudy (anonymous):

O_o lol i don't get it, what's the final answer? :D

OpenStudy (strawberry17):

hehe :D

OpenStudy (anonymous):

11/66?

OpenStudy (strawberry17):

I guess... It sounds like @E.ali know what he/she is talking about! :D

OpenStudy (anonymous):

@strawberry17 : We should find Y Delta . K ?

OpenStudy (anonymous):

@E.ali always right lol :D

OpenStudy (anonymous):

Every body can fall !:)

OpenStudy (strawberry17):

@E.ali what is Y Delta??? :O

OpenStudy (anonymous):

no wait... I don't get the question now.. round to the nearest tenth?

OpenStudy (anonymous):

Means Y s change :)

OpenStudy (anonymous):

@E.ali whats the answer lol?

OpenStudy (strawberry17):

LOL :)

OpenStudy (anonymous):

3.3

OpenStudy (anonymous):

Rounded solution.

OpenStudy (anonymous):

O_o okay, thanks!

OpenStudy (strawberry17):

but we all got different solutions @AnimalAin so how will he know that is right? :)

OpenStudy (anonymous):

@AnimalAin : What answer did you get ?

OpenStudy (anonymous):

10/3. Note my post at the beginning of the thread. It rounds to 3.3.

OpenStudy (anonymous):

What are you saying a bout ??? Is it hard to understand ? I said : 66/11=6 Y s change and then we have : 20 *6=120=x But whats the round to the nearest tenth ??? I cant transelate it ?

OpenStudy (anonymous):

Inversely proportional relations can be shown as xy=c. In this case, based on the given information, xy=220. If we substitute 66 for y, we get 10/3 for x, which rounds to 3.3.

OpenStudy (strawberry17):

Did you get the rounded answer?? @charlieberzak

OpenStudy (anonymous):

3.3

OpenStudy (strawberry17):

that is according to AnimalAin, but you mentioned that E.ali is always right :) so his answer would be 120... so just go with what seems right.... :)

OpenStudy (strawberry17):

it sounds like AnimalAin is right though, cause his answer actually had to be rounded to a tenth.